OpenStudy (anonymous):

Which is the equation of the line, in point-slope form, that has a slope of –4 and passes through the point (7, 1)? (Points : 1) y – 7 = 4(x – 1) y – 1 = –4(x + 7) y – 1 = –4(x – 7) y – 7 = –4(x – 1)

OpenStudy (anonymous):

Two lines are perpendicular if their slopes are opposite recipricals. That means that the equation to compare the slopes is: m1 = -1/m2, where m1 and m2 are the slopes of the two lines. The equation for the line that is perpendicular to the line we want is y =-1/4x -5, which means its slope is -1/4. m1 = -1/(-1/4) Equation for perpendicular slope with -1/4 substituted for m2 m1 = 4 Simplified Now that we have the slope of the line that we want (4) and a point (-1,7), we can use point-slope formula to find the equation of the line that we want. y - y1 = m(x - x1) Point-slope formula y - 7 = 4(x - (-1)) Substitution y - 7 = 4(x + 1) Subtracting a negative means adding a positive y - 7 = 4x + 4 Distributive property y - 7 + 7 = 4x + 4 + 7 Add 7 to each side y = 4x + 11 Simplif

OpenStudy (anonymous):

*simplify